### v2

# Two-photon interference between mutually-detuned resonance fluorescence signals scattered off a semiconductor quantum dot

Abstract
Radiative linewidth of a two-level emitter (TLE) ultimately determines the bandwidth it can offer for quantum information processing. However, no prior experiment has so far been performed to examine the effect of driving detuning on indistinguishability of photons scattered off a TLE, a parameter that is crucial for photonic quantum computing. Here, we perform post-selective two-photon interference experiments between mutually-detuned resonance fluorescence signals from an InAs quantum dot embedded in a micropillar cavity. Our results suggest that indistinguishability among photons scattered off a quantum dot is inherently insensitive to the driving laser’s detuning, as straightforwardly predicted by the resonance fluorescence model that systematically treats all scattered photons as spontaneous emission.
Impinging upon a semi-transparent beam-splitter, two indistinguishable single photons will coalescence into either of the two output ports, and never exit separately. Referred to as Hong-Ou-Mandel (HOM) effect 1, this phenomenon is the foundation for quantum information protocols such as repeater-assisted quantum communication 2 and linear-optic quantum computing 3. Single photons can be generated through exciting an atom-like system 4, 5, 6, which by nature emits one photon at a time. In this regard, semiconductor quantum dots (QDs) are arguably the most attractive platform because they can be integrated into monolithic structures for cavity enhancement 5, 7, 8 and highly indistinguishable single photons 9, 10, 11.
Resonant excitation into the lowest excited state allows a QD system to be simplified as a two-level emitter (TLE) 12. Under a weak coherent pump (Heitler regime), the resulting resonance fluorescence (RF) signal locks to the driving laser’s coherence 13, 14, 15, 16 while exhibiting single-photon characteristic. To a first approximation, the QD appears to behave like a passive and elastic scatterer, which led to an optimistic idea to circumvent environment-induced dephasing inherent in all solid-state systems and to generate single photons with indistinguishability “fundamentally” defined by the excitation laser 14. A question then arises. In the presence of an excitation detuning, will the single photons remain indistinguishable to their counterparts generated under strictly resonant excitation? An affirmative answer would be favorable for quantum information processing because it will then permit arbitrary modulation of the driving laser and the shaping of the RF beam accordingly without loss of photon indistinguishability. However, no experiment has so far been dedicated to explicitly examine this question.
Fig. 1: Dual-color two-photon interference setup. a, High-resolution RF spectra measured with a Fabry-Pèrot interferometer (FPI); The dashed line represents the QD-micropillar reflectance spectrum measured under $\bar{n}=0.05$. b, The auto-correlation function $g^{(2)}(0)$ values measured under $\bar{n}=0.05$ using a Hanbury Brown-Twiss setup (HBT). The red dashed line shows the cavity reflectivity under a strong excitation of 465 nW, corresponding to $\bar{n}=139$. BS: beam splitter; SPD: single-photon detector.
Fig. 2: Dual-color HOM interference results under two different mutual detunings: $\Delta=0$ GHz (left column) and $\Delta=0.5$ GHz (right column). a, b, HOM interference traces for the parallel (blue) and orthogonal (red) polarization settings; c, d, Magnified view of the data shown in panels a and b; e, f, TPI visibility. In all panels, experimental data are shown as dots while theoretical simulations are displayed as solid lines. In this set of measurements, the excitation flux is fixed at $\bar{n}=0.05$.
Here, we subject an InAs QD-micropillar device to dual-color excitation, i.e., from two different wavelengths, using the setup shown in Fig. 1, and evaluate indistinguishability of mutually-detuned RF signals using post-selective two-photon interference (TPI) 17. Two continuous-wave (CW) lasers (not shown) are carved into 0.84 MHz square waves before being temporally interleaved to excite the QD device. Symmetrical detuning is applied: one laser is detuned by $+\frac{\Delta}{2}$ while the other by $-\frac{\Delta}{2}$ from the QD resonance ($\nu_{0}$). The resulting RF tracks the lasers’ frequencies and consists of RF signal pulses of alternating colors. The pulses enter an asymmetric Mach-Zehnder interferometer (AMZI) of 595 ns delay that matches the laser carving frequency, so that the RF signals of different colors are temporally aligned when meeting at the AMZI’s exit beam splitter. The interference results are registered by a pair of superconducting nanowire single photon detectors (SPDs) of 59 ps temporal resolution.
The InAs QD sample was fabricated from a wafer grown by molecular beam epitaxy on a (001) semi-insulating GaAs substrate and embedded into a double distributed-Bragg-reflector (DBR) micropillar cavity with a diameter of 2.4 $\mu$m and a quality factor (Q) of 9350. A confocal microscope (not shown) collects the RF at the same polarization as the excitation, thanks to the low cavity reflection of the device 18, 16. The sample is cooled by a closed-cycle cryostat to 5.6 K to ensure the QD’s resonance with the micropillar cavity of 911.485 nm. Under pulsed resonant excitation, we characterize the QD to have a radiative lifetime ($T_{1}$) of $74$ ps, corresponding to a natural linewidth of $\gamma_{\parallel}:=\frac{1}{2\pi T_{1}}=2.15$ GHz. In all subsequent experiments, we denote the excitation flux upon the device surface using the relation $\bar{n}=P_{\mathrm{in}}T_{1}/h\nu$, where $P_{\mathrm{in}}$ corresponds to the CW incident optical power.
Before the dual-color experiment, we characterize the QD under excitation by a single laser. The excitation flux is kept at $\bar{n}=0.05$, which is conservatively set low to operate in the Heitler regime. Fig. 1a shows a set of high-resolution RF spectra measured by a Fabri-Pèrot interferometer (FPI). At each detuning, the RF peak inherits the laser’s optical frequency and has a full-width at half-maximum width (FWHM) that is limited by the FPI resolution ($20$ MHz). The peak intensity decreases as the detuning widens, which is a result of the reduced cavity enhancement. At the measurement step-size of 0.5 GHz, the adjacent RF spectra are fully distinguished. Fig. 1b shows the corresponding auto-correlation function results, measured with a Hanbury Brown-Twiss (HBT) setup. Across the detuning range, $g^{(2)}(0)$ is consistently lower than 0.5, indicating the non-classical nature of each RF signal. The lowest $g^{(2)}(0)=0.026$ is obtained at 0-detuning. It gradually deteriorates as the detuning increases, and exceeds $0.10$ when the detuning reaches ±$\pm 2$ GHz. We attribute the deterioration in single-photon purity to the increased laser background, as evidenced by the cavity reflectivity (red dashed line), which was measured under strong laser excitation.
We use the setup in Fig. 1 for dual-color excitation experiments. By rotating the half-wave plate ($\lambda/2$) in the long arm of the AMZI, the polarization alignment between the incoming photon beams can be made either parallel or orthogonal. In the latter case, photons are fully distinguishable. We measure the normalized second-order correlation functions under both polarization settings, from which we calculate the TPI visibility ($\mathcal{V}$) using 17
$\mathcal{V(\tau)}=\frac{g^{(2)}_{\perp}(\tau)-g^{(2)}_{\parallel}(\tau)}{g^{(2)}_{\perp}(\tau)},$ (1)
where $\tau$ represents the coincidence time delay while $g^{(2)}_{\perp}(\tau)$ and $g^{(2)}_{\parallel}(\tau)$ are the normalized second-order correlation functions under the orthogonal and parallel polarization settings, respectively.
As a control experiment, we set the two lasers to be nominally identical in wavelength, i.e., $\Delta=0$ GHz, and Fig. 2a shows the measured HOM traces (dots). Here, the coincidence probabilities are normalized to the average rate at long delays ($\tau>200$ ns). With orthogonal polarizations, the photons are fully distinguishable and therefore no interference takes place. Hence, the corresponding $g^{(2)}_{\perp}(\tau)$ trace displays a single, sharp dip with a depth of $\sim$0.5 and a width of about 100 ps, arising from the single-photon nature of the RF streams. When the polarizations are aligned, photons become largely indistinguishable, two-photon coalescence takes place thus resulting in a more pronounced HOM dip at $\tau=0$. Additionally, the correlation function exhibits a broad suppression extending over a timescale of ±$\pm 100$ ns. Over nanosecond timescale (Fig. 2c), $g_{\perp}^{(2)}$ and $g_{\parallel}^{(2)}$ exhibit flat values for $\tau\gg T_{1}$ but with $g_{\parallel}^{(2)}$ being considerably lower than $g_{\perp}^{(2)}$. This is a common observation in the HOM measurements for resonantly excited QDs 16, 15. Using Eq. 1, we obtain the TPI visibility shown in Fig. 2e.
Next, we measure the second-order HOM correlation functions under dual-color excitation, with the result of a $\Delta=0.5$ GHz detuning shown in Fig. 2b. While the $g^{(2)}_{\perp}$ data is qualitatively identical to the non-detuned case, the $g^{(2)}_{\parallel}$ data exhibits symmetrical oscillations around $g^{(2)}_{HOM}=1$. The oscillation amplitude decreases as the delay enlarges, and its lower envelope closely resembles to the non-detuned trace. Under a magnified view (Fig. 2d), the oscillation is discerned to have a period of $1/\Delta$, similarly to the asynchronous TPI result between frequency-offset laser fields 19. As a display of the HOM effect, photon coalescence leads $g^{(2)}_{\parallel}(0)$ to be substantially lower than the oscillation minima. Figure 2f shows the TPI visibility, whose trace has a substantially different appearance from that in Fig. 2e. The $\mathcal{V}$ value oscillates (negative data not shown) with the oscillation maxima upper-bounded to the TPI baseline measured in the non-detuning case. Nevertheless, the $\mathcal{V}$ maximum at $\tau=0$ is about identical to that of the non-detuned case.
To explain the above observations, we resort to the recently introduced RF model 16 that uniquely treats all the RF photons as a result of absorption and re-emission. Considering the TLE and its photon emission together, this RF model adopts a pure-state description for the joint system of the TLE and its emission, although the TLE or the RF emission separately is well known to be in a mixed state. As derived in Appendix A, the quantum state of the light-matter system can be written as
|$\ket{\psi}=\sqrt{p_{0}}\ket{0,g}+\sqrt{p_{1}}{\color[rgb]{0,0,0}e^{-i2\pi\nu t}}\frac{\ket{0,e}+\ket{1,g}}{\sqrt{2}}$ (2)
where the state |$\ket{0}$ (|$\ket{1}$) corresponds to the absence (presence) of a freshly emitted photon, |$\ket{g}$ (|$\ket{e}$) the TLE’s ground (excited) state and $\nu$ the driving laser’s frequency that can be detuned from the TLE’s transition frequency $\nu_{0}$. The terms $p_{0}$ and $p_{1}$ ($p_{0}+p_{1}=1$) represent the probabilities of the TLE system occupying the system ground state |$\ket{0,g}$ and the joint excited state (|$\left(\ket{0,e}+\ket{1,g}\right)/\sqrt{2}$, respectively. The latter is in a superposition form conveying the insight that at any given instant of time, it is not possible to know whether the TLE stays in its excited state |$\ket{e}$ or has decayed to the ground state |$\ket{g}$ and emitted a photon 16. Naturally, |$\ket{1}$ represents a photon emitted by the TLE, implying that the property of the photons is independent of excitation detuning. In the absence of pure dephasing, we have
$p_{0}=\frac{1}{\left(1+2\eta_{ab}\bar{n}\right)},$ (3)
where $\eta_{ab}$ is the TLE’s absorption efficiency at weak pump limit and depends on the excitation detuning.
Eq. 2 allows us to infer the RF’s first-order coherence to be $g^{(1)}(\tau)=p_{0}e^{-i2\pi\nu\tau}$ for $\tau\gg T_{1}$, via evaluation of the interference outcome between the pure quantum state with a delayed copy of itself 16. Through Fourier transform of $g^{(1)}(\tau)$, we readily obtain that the detuned RF spectrum contains a laser-like component, with a spectral weight of $p_{0}$, that inherits the driving laser’s frequency ($\nu$) and coherence time ($T_{L}$). At weak pump limit, the fraction of the laser-like component can be arbitrarily close to unity. Hence, the vast RF spectral distinguishability shown in Fig. 1b can be attributed to the interference of the phase-locked temporal modes of single photons with properties identically defined by the QD.
Fig. 3: Visualization of two-photon coincidence contributions. a, Two mutually-detuned RF beams, with global phases of $\phi_{a}$ and $\phi_{b}$, respectively, meet at a 50/50 beam splitter, and their interference is recorded by two single-photon detectors; b, Four possible cases involving two input photons may contribute to a coincidence event with the two detectors clicking respectively at time $t_{1}$ and $t_{2}=t_{1}+\tau$. Cases (ii) and (iii), in which each input arm contributes one photon, share the common global phase ($\phi_{a}+\phi_{b}$) and therefore their quantum probability amplitudes interfere. Cases that involve three or more input photons exist and are not shown.
The dual-color HOM data can be elegantly explained through the schematic shown in Fig. 3. Without loss of generality, we fix the entry of the red (blue) detuned RF signal to the top (bottom) input of the 50/50 beam splitter. Their global phases are respectively depicted as $\phi_{a}$ and $\phi_{b}$, and are not mutually locked. A coincidence of photon clicks at $t_{1}$ and $t_{2}=t_{1}+\tau$ by two detectors can arise from four possible cases: both photons came from the same entry (i and iv) or they entered from different entries (ii and iii). Here, we temporally ignore cases involving three or more input photons to avoid complication. Among the four possible cases, cases (ii) and (iii) share the same global phase ($\phi_{a}+\phi_{b}$) and therefore interference occurs after their transformation through the beam splitter. The phase difference is $2\pi\Delta\tau+\pi$, where $2\pi\Delta\tau$ arises from the RF detuning and an extra $\pi$ phase gains from the beam splitter. The interference will cause the total coincidence probability to oscillate around an intermediate value that is defined by cases (i) and (iv).
In the single-photon limit that each input arm contributes at most one photon between time bins $t$ and $t+\tau$, i.e., $g^{(2)}(\tau)\to 0$, the coincidence probability amplitudes from cases (ii) and (iii) can cancel completely when $\Delta\cdot\tau$ is an integer (due to a $\pi$ phase shift from the beam splitter), but interfere constructively when $\Delta\cdot\tau$ becomes an half integer. The HOM correlation function is phase-dependent:
$g^{(2)}_{HOM}(\tau)=\frac{1-\mathcal{M}\cos{2\pi\Delta\tau}}{2},$ (4)
where $\mathcal{M}=|\langle 1_{+}|1_{-}\rangle|^{2}$ represents indistinguishability between temporally-aligned single photon states (|$\ket{1_{\pm}}$) generated under opposite detunings. We remark that states |$\ket{1_{\pm}}$ represent spontaneous emission and therefore their indistinguishability is affected by the QD environment fluctuation but independent of the excitation detuning. $\mathcal{M}$ can also be adjusted through the half-wave plate in the experimental setup (Fig. 1). With orthogonal polarizations, $\mathcal{M}=0$ and $g^{(2)}_{\perp}(0)$ has a theoretical minimum of 0.5.
When the coincidence delay is much greater than the QD’s exciton radiative lifetime, we arrive at the other limit that $g^{(2)}(\tau)=1$ and the probability of time bin $t+\tau$ containing a photon is independent of that at time bin $t$. Notably, by using the pure-state of Eq. 2, we can rigorously derive the second-order HOM correlation function for $\tau\gg T_{1}$ as
$g_{HOM}^{(2)}(\tau)=\frac{1}{2}+\frac{1}{2}\left(1-\mathcal{M}p_{0}^{2}\cos{2\pi\Delta\tau}\right).$ (5)
The first term $\frac{1}{2}$ arises from cases (i) and (iv), i.e., both photons within a coincidence event came from the same beam splitter input. The second term represents the interference between the quantum probability amplitudes of cases (ii) and (iii), but which can no longer cancel completely because of the non-negligible probability of each input arm contributing two photons. The detailed derivation of Eq. 5 is presented in Appendix B.
To model the situation of an arbitrary coincidence delay, we bridge the above two limits using the auto-correlation function $g^{(2)}(\tau)$ and obtain
$g_{HOM}^{(2)}(\tau)=$ $g^{(2)}(\tau)\times\left(1-\frac{1}{2}\mathcal{M}p_{0}^{2}\cos{2\pi\Delta\tau}\right)$ (6)
$+(1-g^{(2)}(\tau))\times\frac{1-\mathcal{M}\cos{2\pi\Delta\tau}}{2}.$
As check, one has that for $\tau\gg T_{1}$, Eq. 6 reduces to the form of Eq. 5 using $g^{(2)}(\tau)=1$. At the other end for $\tau=0$, ideal single-photon purity of $g^{(2)}(0)=0$ leads the equation to standard post-selective HOM interference, i.e., Eq. 4.
We simulate the HOM experiments using Eq. 6, with results shown as solid lines in Fig. 2. The auto-correlation function is directly taken from the experimental HBT data using $g^{(2)}(\tau)=\sqrt{g_{+}^{(2)}(\tau)g_{-}^{(2)}(\tau)}$, where $g_{\pm}^{(2)}(\tau)$ are the measured auto-correlation functions under the corresponding positive ($+\frac{\Delta}{2}$) or negative ($-\frac{\Delta}{2}$) excitation detuning. We use $\mathcal{M}=0.94$ ($\mathcal{M}=0$) for the parallel (orthogonal) polarization setting. We also take into account the finite coherence time between the driving lasers, which deteriorates the TPI visibility at large delays. As reported in Figs. 2ad, the simulation faithfully reproduces all the measurement data for both polarization settings, both excitation detunings and all delays.
In the above simulation, we choose $p_{0}=0.92$ corresponding to an excitation flux of $\bar{n}=0.05$. As $p_{0}$ decreases with a strengthening driving intensity, we expect the oscillation amplitude of the detuned HOM correlation to dampen according to Eq. 5. To test, we measure power-dependent dual-color HOM traces with the results (symbols) shown in Fig. 4a. Fitting the correlation oscillations using Eq. 5 (solid lines), we extract the oscillation amplitude ($\frac{1}{2}\mathcal{M}p_{0}^{2}$) and plot its dependence on the excitation flux in Fig. 4b. The measured dependence (symbol) is in agreement with the theoretical simulation (dashed line) using Eq. 3.
Fig. 5 reports the measured 0-delay HOM correlation values as a function of excitation detuning. We observe a gradual increase of $g^{(2)}_{\parallel}(0)$, which is expected from the RF’s deteriorating single-photon purity at large excitation detunings (see Fig. 1b). However, $g^{(2)}_{\perp}(0)$ exhibits an opposite trend, and unexpectedly decreases as the excitation detuning exceeds 0.5 GHz. At $\Delta=4$ GHz, $g^{(2)}_{\perp}(0)$ is measured to be just 0.31, which is significantly below the theoretical minimum of 0.5 for distinguishable photons. Often, $g^{(2)}_{\perp}(0)<0.5$ can be caused by an intensity imbalanced AMZI, which is a common flaw in a HOM setup 15. For example, at the extreme condition with one AMZI arm completely blocked, the HOM interferometer becomes equivalent to a HBT setup. We have ruled out the AMZI as the possible cause for $g^{(2)}_{\perp}(0)<0.5$ in our experiment. Consequently, $g^{(2)}_{\perp}(0)<0.5$ can be caused only by the QD source and is probably related to the large excitation detuning. While the exact mechanism is yet to be pinned down, it is likely to have a similar effect on the parallel polarization setting and hence Eq. 1 can remain meaningful for calculating the TPI visibility.
The experimental TPI visilibity is shown as solid squares in Fig. 5. When $\Delta>0.5$ GHz, $\mathcal{V}(0)$ decreases monotonically with the excitation detuning. The TPI visibility $\mathcal{V}(0)$ is closely linked to photon indistinguishability. In an ideal scenario of $g^{(2)}(0)=0$, it is straightforward to deduce $\mathcal{V}(0)=\mathcal{M}$ from Eqs. 1 and 6. As all RF photons are treated as spontaneous emission, the pure-state RF model (Eq. 2) immediately predicts their indistinguishability to be independent of excitation detuning. Apparently, the measured $\mathcal{V}(0)$ does not follow the predicted behavior for photon distinguishability. A number of factors can contribute to this discrepancy. Firstly, single-photon purity degrades at large detunings due to the increasing laser background. Secondly, the detectors’ time resolution prevents from resolving the HOM interference dip, especially when the dip is modulated by a large detuning frequency. Finally, the measured $g^{(2)}_{\perp}(0)<0.5$ at large detunings effectively reduce the measured $\mathcal{V}(0)$ value. By taking into account these factors, we construct a model (see Appendix D) and calculate the post-selective TPI visibility with results as the red line in Fig. 5. Here, each required $g^{(2)}(0)$ value is interpolated from the experimental measurement shown in Fig. 1b. In accordance with the RF model (Eq. 2), we use a fixed value of $\mathcal{M}=0.94$, which was conservatively chosen based on the first-order coherence measurement shown in Appendix C. The result (green line) is in excellent agreement with the experimental data (squares), thereby supporting the RF model based on the pure-state description.
Fig. 4: Excitation flux dependence of the HOM correlation under the parallel polarization setting. The excitation detuning is fixed at $\Delta=0.5$ GHz. a, HOM correlation traces. Symbols (solid lines) denote experimental data (theoretical fits). b, The oscillation amplitude (symbols) extracted from the data shown in panel a as well as the fit (dashed line) using Eq. 3.
Fig. 5: Excitation detuning dependencies of the 0-delay TPI visibility and the HOM correlations. For all measurements, the excitation flux is fixed at $\bar{n}=0.05$.
To conclude, we have experimentally investigated post-selective TPI between mutually detuned RF signals scattered off a single QD. By invoking the model that treats all the RF photons as result of spontaneous emission (Eq. 2), it is possible to find full correspondence between the predictions of the model and the experimental results. Beyond validating the theoretical model, these findings are also highly relevant to quantum information processing, as we demonstrated that indistinguishable photons can be generated by an emitter, such as a quantum dot, irrespective of the excitation detuning.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grants 62250710162, 12494600, 12494604 and Beijing Natural Science Foundation under grant IS23011.
Appendix A Model for a coherently-driven two-level system
Fig. 6: Alternative pictures for a two-level emitter under weak, detuned coherent driving. a, Elastic/passive scattering; b, Spontaneous emission into temporal modes that are phase-locked by the excitation laser. $\nu_{0}$: Transition gap between the ground (|$\ket{g}$) and exited (|$\ket{e}$) states; $\Delta$: Excitation detuning.
As depicted in Fig. 6a, a TLE under a weak coherent excitation is usually considered as a passive scatterer 20, as the scattered light inherits the driving laser’s energy and linewidth (coherence). On the other hand, the scattered signal is known to exhibit photon anti-bunching over the time scale of the TLE’s radiative lifetime, which can be several orders of magnitude shorter than the driving laser’s coherence time. This sharp contrast led to the term of “sub-natural” 21 linewidth, and the reconcilliation could be understood, with careful elaboration, in terms of the interference between the passively scattered photons and the actively emitted photons resulting from actual photon absorption and re-emission 22.
We have recently proposed a simple model 16 that treats all RF photons systematically as an active result of absorption and re-emission. This model successfully accounts for both the RF’s first-order coherence and photon anti-bunching under strict resonant condition. Extending to a detuned excitation, it immediately predicts that all RF photons from the same TLE remain identical to one another irrespective to the excitation detuning, even when the RF fields may exhibit vastly different central frequencies. Fig. 6b schematically summarizes the RF model under a detuned excitation: the properties of the single photons are defined by the two-level emitter but the temporal modes they radiate into are phase-locked by the excitation laser.
For strict resonant excitation, we have previously derived the pure-state description of Eq. 2 when considering the TLE and its photon emission jointly 16, although the TLE and the emission separately are mixed states. Following similar derivation steps, here we generalize the pure-state description to cover the presence of excitation detuning. Our starting point is the model 23 of a two-level system in the frame that rotates with the driving laser and in the CW regime, i.e.,
$H=-\frac{\Omega}{2}\left(e^{-i\phi}\sigma^{\dagger}+e^{i\phi}\sigma\right)+\frac{\Delta}{2}\sigma_{z},$ (7)
where $h=1$, $\sigma=|g\rangle\langle e|$ ($\sigma^{\dagger}=|e\rangle\langle g|$) is the atom lowering (raising) operator, $\sigma_{z}=|g\rangle\langle g|-|e\rangle\langle e|$ is the Pauli $z$ matrix, $\Omega$ is the Rabi frequency and $\Delta:=\nu-\nu_{0}$ is the excitation detuning. $\phi$ is the phase of the dipole matrix element $\mathbf{p}_{ge}=|\mathbf{p}_{ge}|e^{i\phi}$. Since it does not produce an observable effect, we let $\phi=\pi/2$.
If we consider the spontaneous emission with the Markov approximation, we have
$\frac{d}{dt}\rho=\mathcal{L}\rho=-i[H,\rho]+{\color[rgb]{0,0,0}\mathcal{D}_{\gamma_{\parallel},\sigma}[\rho_{a}]+\mathcal{D}_{\gamma^{*}/2,\sigma_{z}}[\rho_{a}]},$ (8)
with $\gamma_{\parallel}$ being the spontaneous emission rate, $\gamma^{*}$ the pure dephasing rate of the emitter, $\rho_{a}$ is the density matrix of a two-level atom, i.e., TLE and $\mathcal{D}$ the dissipation superoperator given by
$\mathcal{D}_{x,A}[\rho_{a}]=x\left[A\rho A^{\dagger}-\frac{1}{2}(A^{\dagger}A\rho+\rho A^{\dagger}A)\right].$ (9)
From an initial state $\rho(0)=|g\rangle\langle g|$, we analytically solve the system state at time $t$ by
$\rho(t)=e^{t\mathcal{L}}\rho(0).$ (10)
Since we are only interested in the system at equilibrium, we can simplify our analysis by considering the steady-state value as $t\rightarrow\infty$ to get
$\rho_{gg}(\infty)=\frac{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+4\Delta^{2}\right)+\gamma_{\perp}\Omega^{2}}{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+4\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}},$ (11)
$\rho_{ee}(\infty)=\frac{\gamma_{\perp}\Omega^{2}}{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}},$ (12)
$\rho_{ge}(\infty)=\rho_{eg}^{*}(\infty)=\frac{\gamma_{\parallel}\left(\gamma_{\perp}+i\Delta\right)\Omega}{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}},$ (13)
where $\gamma_{\perp}=\gamma_{\parallel}/2+\gamma^{*}$ is the spectrally-broadened emitter linewidth. Then, the steady-state density matrix of the TLE alone can be written as:
[…]
$\left.+\gamma_{\|}\left(\gamma_{\perp}-i\Delta\right)\Omega|e\rangle\langle g|+\gamma_{\perp}\Omega^{2}|e\rangle\langle e|\right\}.$
With the hypothesis that the driving laser is perfectly filtered from the radiation emitted by the source, we can leverage the relation of proportionality between the source dipole and the field operator of the collected light, i.e., $a{\color[rgb]{0,0,0}\propto}\sigma$. Therefore, we might as well let $a=B\sigma$, where $B$ is a proportional coefficient. In this way, we always have that the state of light normalized by the intensity corresponds to the steady-state of the emitter.
Formally, we can perform a tomographic light analysis by examining its coherence and intensity at a specific instant $t$, once the emitter has reached a steady state:
$\langle a(t)\rangle_{\infty}=\operatorname{Tr}\left[B\sigma\rho_{\infty}\right]=B\rho_{eg}(\infty)$ (15)
$\left\langle a^{\dagger}(t)a(t)\right\rangle_{\infty}=\operatorname{Tr}\left[B^{2}\sigma^{\dagger}\sigma\rho_{\infty}\right]=B^{2}\rho_{ee}(\infty).$ (16)
The first order coherence is then obtained by normalizing the squared magnitude of the coherence by the intensity:
$|g_{\infty}^{(1)}|=\frac{|B\rho_{eg}(\infty)|^{2}}{B^{2}\rho_{ee}(\infty)}=\frac{\gamma_{\parallel}^{2}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)}{2\gamma_{\perp}\left[\gamma_{\parallel}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+\gamma_{\perp}\Omega^{2}\right]}.$ (17)
It can be seen that the specific value of $B$ does not affect $g^{(1)}_{\infty}$ of fluorescence. Since a single excitation of the TLE can be converted into a single photon, thus the fluorescence density matrix at the steady state can be determined as well,
[…]
$\left.+\gamma_{\|}\left(\gamma_{\perp}-i\Delta\right)\Omega|1\rangle\langle 0|+\gamma_{\perp}\Omega^{2}|1\rangle\langle 1|\right\}.$
After obtaining the reduced density matrices of the TLE and fluorescence individually, we can infer the density matrix of the TLE-photon joint system. Under the equilibrium condition of continuous-wave laser driving, at any given time $t$, the density matrix state vector of the joint system of the two-level atom and fluorescence only needs to consider $|g0\rangle$, $|e0\rangle$, and $|g1\rangle$, while excluding higher-order terms such as $|e1\rangle$ and $|g2\rangle$. This is primarily based on exciting experimental observations and theoretical results, where $g^{(2)}(0)=0$ is always valid, regardless of whether the excitation is strong or weak. Therefore, the instantaneous atom-photon density matrix can be written in the form
$\rho_{ap}=\left(\begin{array}[]{lll}\rho_{g0,g0}&\rho_{g0,g1}&\rho_{g0,e0}\\ \rho_{g1,g0}&\rho_{g1,g1}&\rho_{g1,e0}\\ \rho_{e0,g0}&\rho_{e0,g1}&\rho_{e0,e0}\end{array}\right).$ (19)
In the absence of pure dephasing, i.e., $\gamma^{*}=0$, we obtain the pure-state form of Eq. 2 and
$p_{0}=\frac{{\color[rgb]{0,0,0}\gamma_{\parallel}}^{2}+4\Delta^{2}}{{\color[rgb]{0,0,0}\gamma_{\parallel}}^{2}+4\Delta^{2}+2\Omega^{2}}=\frac{1}{1+2\eta_{ab}\bar{n}},$ (20)
$p_{1}=1-p_{0}=\frac{2\eta_{ab}\bar{n}}{1+2\eta_{ab}\bar{n}},$ (21)
where $\frac{\Omega^{2}}{{\color[rgb]{0,0,0}\gamma_{\parallel}}^{2}+4\Delta^{2}}=\eta_{ab}\bar{n}$ and $\eta_{ab}$ represents the photon absorption efficiency at weak pump limit.
Appendix B Derivation of $g^{(2)}_{\parallel}(\tau)$ under steady-state condition ($\tau\gg T_{1}$)
In this section, we will discuss the coincidence counts for different incident configurations depicted in Fig. 3 from the perspective of the input photon states. We start from the pure-state description of Eq. 2. Tracing out the TLE component state (|$\ket{g}$, |$\ket{e}$), the remaining photon sub-system is in a mixed state, which is equivalent to the pure photon state |$\ket{\psi}_{ph}=\sqrt{p_{0}}\ket{0}+\sqrt{p_{1}}e^{i2\pi\nu t}\ket{1}$ attenuated by a channel of 50% loss. Conveniently, channel losses do not affect the analysis of the first-order or second-order correlation functions as we have previously proved 16, so we are able to use the pure photon state in the subsequent derivation.
We consider two mutually-detuned RF signals entering a 50/50 beam splitter via ports $a$ and $b$, and exiting from ports $c$ and $d$. Under steady-state condition ($\tau\gg T_{1}$), the input state can be expressed as the product of photon states at different instants of time for ports $a$ and $b$ as $|\psi_{in}\rangle=|\psi\rangle_{a}|\psi\rangle_{b}$, where:
$|\psi\rangle_{a}=|\psi\rangle_{a,t}|\psi\rangle_{a,t+\tau}=\left(\sqrt{p_{0}}|0\rangle_{a,t}+\sqrt{p_{1}}|1\rangle_{a,t}e^{-i2\pi\nu_{a}t}\right)\left(\sqrt{p_{0}}|0\rangle_{a,t+\tau}+\sqrt{p_{1}}|1\rangle_{a,t+\tau}e^{-i2\pi\nu_{a}(t+\tau)}\right),$ (22)
$|\psi\rangle_{b}=|\psi\rangle_{b,t}|\psi\rangle_{b,t+\tau}=\left(\sqrt{p_{0}}|0\rangle_{b,t}+\sqrt{p_{1}}|1\rangle_{b,t}e^{-i2\pi\nu_{b}t}\right)\left(\sqrt{p_{0}}|0\rangle_{b,t+\tau}+\sqrt{p_{1}}|1\rangle_{b,t+\tau}e^{-i2\pi\nu_{b}(t+\tau)}\right).$
In the above formula, $\nu_{a}=\nu_{0}-\frac{\Delta}{2};$ $\nu_{b}=\nu_{0}+\frac{\Delta}{2}$, $\Delta$ is the frequency detuning relative to the center frequency $\nu_{0}$ and $\tau$ is the coincidence time interval. According to the transformation relation of a beam splitter, the input optical field annihilation operator can be given by
$c_{t}d_{t+\tau}$ $=\frac{1}{\sqrt{2}}(a_{t}+b_{t})\times\frac{1}{\sqrt{2}}(a_{t+\tau}-b_{t+\tau})$ (23)
$=\frac{1}{2}(a_{t}a_{t+\tau}-a_{t}b_{t+\tau}+a_{t+\tau}b_{t}-b_{t}b_{t+\tau}),$
where $a_{t}$ and $a_{t+\tau}$ act on states $|\psi\rangle_{a,t}$ and $|\psi\rangle_{a,t+\tau}$. Similarly, the remaining annihilation operators adhere to this convention. Then, the count probabilities of detectors in ports $c$ and $d$ can be written as
$P(\tau)$ $=\langle\psi_{in}|c_{t}^{\dagger}d_{t+\tau}^{\dagger}c_{t}d_{t+\tau}|\psi_{in}\rangle$ (24)
$=\left|\frac{1}{2}a_{t}a_{t+\tau}|\psi_{in}\rangle\right|^{2}+\left|-\frac{1}{2}b_{t}b_{t+\tau}|\psi_{in}\rangle\right|^{2}+\left|-\frac{1}{2}a_{t}b_{t+\tau}|\psi_{in}\rangle+\frac{1}{2}a_{t+\tau}b_{t}|\psi_{in}\rangle\right|^{2}$
$=\left|\frac{1}{2}p_{1}e^{-i2\pi\nu_{a}(2t+\tau)}|00\rangle_{a}|\psi\rangle_{b}\right|^{2}+\left|-\frac{1}{2}p_{1}e^{-i2\pi\nu_{b}(2t+\tau)}|00\rangle_{b}|\psi\rangle_{a}\right|^{2}$
$+\left|-\frac{1}{2}(p_{0}p_{1}e^{-i2\pi 2\nu t}e^{-i2\pi 2\nu_{b}\tau}|0000\rangle+p_{1}\sqrt{p_{0}p_{1}}e^{-i2\pi 2\nu t}e^{-i2\pi 2\nu_{b}(t+\tau)}|0010\rangle+p_{1}\sqrt{p_{0}p_{1}}e^{-i2\pi 2\nu(t+\tau)}e^{-i2\pi 2\nu_{a}t}|0100\rangle\right.$
$\left.+p_{1}{}^{2}e^{-i2\pi 2\nu(2t+\tau)}|0110\rangle)+\frac{1}{2}(p_{0}p_{1}e^{-i2\pi 2\nu t}e^{-i2\pi 2\nu_{a}\tau}|0000\rangle+p_{1}\sqrt{p_{0}p_{1}}e^{-i2\pi 2\nu(t+\tau)}e^{-i2\pi 2\nu_{b}\tau}|0001\rangle\right.$
$\left.+p_{1}\sqrt{p_{0}p_{1}}e^{-i2\pi 2\nu t}e^{-i2\pi 2\nu_{a}(t+\tau)}|1000\rangle+{p_{1}}^{2}e^{-i2\pi 2\nu(2t+\tau)}|1001\rangle)\right|^{2}$
$=\frac{1}{4}p_{1}^{2}+\frac{1}{4}p_{1}^{2}+\left|\frac{1}{2}(p_{0}p_{1}e^{-i2\pi 2\nu t}(e^{-i2\pi 2\nu_{a}\tau}-e^{-i2\pi 2\nu_{b}\tau})|0000\rangle\right|^{2}+\frac{1}{4}\times 2(p_{1}^{3}p_{0}+p_{1}^{3}p_{0}+p_{1}^{4})$
$=\frac{1}{4}p_{1}^{2}+\frac{1}{4}p_{1}^{2}+\frac{1}{2}p_{1}^{2}\left[1-p_{0}^{2}\cos{(2\pi\Delta\tau)}\right],$
where the form of photon states (e.g., |$\ket{0000}$) follows the convention that the first two digits correspond to the $t$ and $t+\tau$ time instants of port $a$ respectively, and the last two digits correspond to the $t$ and $t+\tau$ instants of port $b$. Within the final equation, the first and second terms correspond to cases (i) and (iv) illustrated in Fig. 3, while the last term accounts for the interference between the two indistinguishable paths as shown in Fig. 3 cases (ii) and (iii). Considering finite photon indistinguishability, we obtain the second-order HOM correlation function by normalizing to the coincidence probability of $p_{1}^{2}$ at $\tau=\infty$:
$g_{HOM}^{(2)}\left(\Delta\right)=1-\frac{1}{2}\mathcal{M}p_{0}^{2}\cos 2\pi\Delta\tau,~{}|\tau|\gg T_{1},$ (25)
which is identical to Eq. 5 in Main Text.
Fig. 7: Measurement of the RF’s first-order coherence. a, Purpose-built AMZI with a differential delay of 5 ns; b, Normalized count rates at ports $c$ (black) and $d$ (red) under a single laser excitation of $\Delta$ = 0 GHz and $\bar{n}$ = 0.002. The phase drift is caused by varying the temperature of the AMZI.
Appendix C The RF’s first-order correlation function
Photon indistinguishability $\mathcal{M}$ is an important parameter in simulating the HOM correlation traces shown in Fig. 2 as well as the detuning-dependent TPI visibility in Fig. 5. We have chosen a fixed value of $\mathcal{M}=0.94$, and below we justify this choice through measuring the RF’s first-order correlation function $g^{(1)}$.
In measuring $g^{(1)}$, we use a second, purpose-built AMZI with a differential delay of $\tau=5$ ns as shown in Fig. 7a. This delay is much greater than the QD relaxation time but much shorter than the laser coherence time. Under this condition of $T_{1}\ll\tau\ll T_{L}$, the RF’s first-order correlation function can be written as 24, 16
$|g^{(1)}(\tau)|=\sqrt{\mathcal{M}}p_{0},$ (26)
from which we can bound photon indistinguishability to $\mathcal{M}\leq|g^{(1)}|^{2}$ using inequality $p_{0}<1$. Fig. 7b shows the interference fringes, from which we extract the first-order correlation function value of $|g^{(1)}|=0.991\pm 0.008$. This result places an upper bound of 0.98 for $\mathcal{M}$.
Appendix D Modeling the 0-delay HOM results
Here, we describe the calculation method that we use to fit the experimental data of the excitation detuning dependence of the 0-delay TPI visibility $\mathcal{V}(0)$ shown in Fig. 5.
Each experimental $\mathcal{V}(0)$ value is calculated from the corresponding HOM measurements of $g^{(2)}_{\perp}(0)$ and $g^{(2)}_{\parallel}(0)$ using to Eq. 1. However, the measured $g^{(2)}_{\perp}(0)$ values at large detunings fall below the theoretical minimum of 0.5 for completely distinguishable but intensity-balanced RF inputs. To model this irregularity ($g^{(2)}_{\perp}<0.5$), we introduce an empirical parameter $x$ ($0\leq x\leq 1$) to quantify the intensity-imbalance between the RF streams. We fit the measured HOM correlation data using
$g^{(2)}_{HOM}(\tau)=\left[x\times g^{(2)}(\tau)+(1-x)\times g\prime^{(2)}_{HOM}(\tau)\right]\otimes R(\tau),$ (27)
where $g\prime^{(2)}_{HOM}$ represents the HOM correlation function as described by Eq. 6 for always balanced RF signals and the inequality $g\prime^{(2)}_{\perp}(0)\geq 0.5$ holds. $R(\tau)$ represents the detector time response.
Among simulation parameters, $x$ is a fitting parameter and excitation detuning dependent. The auto-correlation functions $g^{(2)}(\tau)$ used are obtained from the de-convolution of the measured HBT data with the detector response $R(\tau)$. We use fixed values of $\mathcal{M}_{\perp}=0$ and $\mathcal{M}_{\parallel}=0.94$ for all excitation fluxes. $R(\tau)$ is a Gaussian function of 59 ps FWHM.